OpenStudy (anonymous):

what is the surface area of a cube whose edges are 16 inches long?

OpenStudy (saifoo.khan):

use the formula: \[\large 6a^2\]

OpenStudy (anonymous):

SA = 6a^2 where a=16 SA=6*16*16 =1536 inches^2